OpenStudy (anonymous):

How to find the first derivative of f(t)=ln(t)/t^5 ?

OpenStudy (campbell_st):

use the quotient rule \[y = \frac{f(x)}{g(x)} ....\frac{dy}{dx} = \frac{g(x)f'(x) - f(x)g'(x)}{(f(x))^2}\] let f(t) = ln(t) then f'(t) = 1/t g(t) = t^5 g'(t) = 5t^4 hope this helps

OpenStudy (anonymous):

im having some trouble simplifying

OpenStudy (campbell_st):

use the drawing tool to show what you have so far

OpenStudy (anonymous):



OpenStudy (campbell_st):



OpenStudy (anonymous):

i got x=0 and x=e^(1/5) for critical numbers

OpenStudy (campbell_st):

well it looks like perhaps a different question... to the one you has asked...

OpenStudy (anonymous):

youre right i forgot to say that i wanted to find the critical numbers with the derivative

OpenStudy (campbell_st):

well then the 1st derivative simplifies to \[f'(t) = \frac{t^4(1 - 5\ln(t))}{t^{10}}\] or \[f(t) = \frac{ 1 - 5\ln(t)}{t^6}\] so to find the critical numbers... 1st you need to realise that t cannot equal 0... as a zero denominator is undefined. So a vertical asymptote exists at t = 0 the next critical number will be when 1 - 5ln(t) = 0 now you need to solve this to find so there is only 1 stationary point \[t = e^{\frac{1}{5}}\] hope this helps

OpenStudy (anonymous):

yup the only critical number is at \[t=e^\frac{ 1 }{ 5 }\]